# The Gestalt Computational Model

### v1

@@ Discussion
We have developed a computational model for Gestalt theory using persistent homology, which integrates Gestalt theory from cognitive psychology with persistent homology in computational topology for the first time. This provides a clear description of the relationship between the global and local aspects in Gestalt theory. Additionally, the capacity of persistent homology to extract global topological features makes it a valuable tool for calculating Gestalt theory. Furthermore, our model has the capacity to combine Gestalt theory with computer vision and artificial intelligence, serving as the computational foundation for visual perception implemented in artificial intelligence.
@@ Methods
Figure 6: Illustration of persistence diagram. a, An example of persistence diagram. As the label shows, the red points in the persistence diagram represent the zero-dimensional persistence diagram, and the blue points represent the one-dimensional persistence diagram. b, The process of clustering points in the given $n$-PD based on persistence.
Data availability
The authors declare that the data supporting the findings of this study are available within the paper and supplementary information files.
@@ Acknowledgments
Y.C. and H.L. carried out the theoretical analysis and wrote the manuscript. Y.C. and J.Y. devised the experiments. All authors participated in editing the manuscript.

### v2

@@ Discussion
We have developed a computational model for Gestalt theory using persistent homology, which integrates Gestalt theory from cognitive psychology with persistent homology in computational topology for the first time. This provides a clear description of the relationship between the global and local aspects in Gestalt theory. Additionally, the capacity of persistent homology to extract global topological features makes it a valuable tool for calculating Gestalt theory. Furthermore, our model has the potential to advance research in the quantitative aspects of Gestalt theory and to serve as a novel method for investigating human visual perception in computational psychology.
@@ Methods
Figure 9: Illustration of persistence diagram. a: An example of persistence diagram. As the label shows, the red points in the persistence diagram represent the zero-dimensional persistence diagram, and the blue points represent the one-dimensional persistence diagram. b: The process of clustering points in the given $n$-PD based on persistence.
@@ Acknowledgments
Y.C. and H.L. carried out the theoretical analysis and wrote the manuscript. Y.C., H.L. and J.Y. devised the experiments. Y.C. and H.L implemented the code of the computational model. All authors participated in editing the manuscript.